Consider the basis S = {v1, v2, v3} for R3, where v1 = (1, 2, 1), v2 = (2, 9, 0), and v3 = (3, 3, 4), and let T:R3 rightarrow R2 be the linear operator such that T(V1) = (1, 0), T(v2) = (- 1, 1), T(v3) = (0, 1) Find a formula for T(x1, x2, x3), and use that formula to find
1
Expert's answer
2020-02-13T09:23:32-0500

In case we want to transform R3 into R2 we need a Matrix of size 3x2."\begin{pmatrix}
 x1 & x4 \\
 x2 & x5 \\
 x3 & x6
\end{pmatrix}" The Matrix gives us 6 variables which we can find by solving 6 corresponding equations 

"\begin{cases}
 x1+2*x2+x3=1\\
x4+2*x5+x6=0\\
2*x1+9*x2=-1\\
2*x4+9*x5=1\\
3*x1+3*x2+4*x3=0\\
3*x4+3*x5+4*x6=1\\
\end{cases}" "\implies" "\begin{cases}
 x1+2*x2+x3=1\\
3*x1+3*x2+4*x3=0\\
2*x1+9*x2=-1\\
\end{cases}
\begin{cases}
2*x4+9*x5=1\\
3*x4+3*x5+4*x6=1\\
x4+2*x5+x6=0\\
\end{cases}" "\implies"

"\begin{cases}
 x1+5*x2=4\\
3*x1+3*x2+4*x3=0\\
2*x1+9*x2=-1\\
\end{cases}
\begin{cases}
2*x4+9*x5=1\\
3*x4+3*x5+4*x6=1\\
x4+5*x5=-1\\
\end{cases}" "\implies"

"\begin{cases}
 x2=9\\
 x1=-41\\
 x3=24\\
\end{cases}
\begin{cases}
x5=-3\\
x4=14\\
x6=-8\\
\end{cases}"

Solution is "\begin{pmatrix}
 -41 & 14 \\
 9 & -3 \\
 24 & -8
\end{pmatrix}".

Then we can get the formula, which is T(x,y,z)= (-41*x+9*y+24*z, 14*x-3*y-8*z)